OpenStudy (anonymous):

can anyone help me with a shell method problem for finding a volume , where y=(x^2)-2and y=-|x| and it revolves around the x-axis

OpenStudy (agent0smith):

The region bound by the two? I'm guessing it's asking for that, see picture. Lemme decide which method would be easiest...

OpenStudy (agent0smith):

Wait, this question is 2 months old... wtf, how did i find it...

OpenStudy (anonymous):

@agent0smith : you can do it for others see. hehehe

OpenStudy (agent0smith):

Yeah, it's not too difficult. You could use the washer method, might be easiest... http://media.wiley.com/Lux/07/39807.nce019.gif Outer radius f(x) = (x^2)-2 and inner radius g(x) = x, then just integrate from where f(x) = g(x), to zero (for simplicity) and double it. \[V = 2 \pi \int\limits_{-1}^{0} \left[ (x^2 -2)^2 - x^2 \right]dx \]

OpenStudy (agent0smith):

Missed that it said to use shell method... well the question is 2 months old anyway...

OpenStudy (agent0smith):

use this for shell method http://media.wiley.com/Lux/17/39817.nce024.gif you'd prob have to use two integrals, one for the x^2 - 2 and one for the -|x|. I might do it later...